OpenStudy (christos):

integral of 1/(u^3) ?

OpenStudy (christos):

@zzr0ck3r u there?

OpenStudy (jhannybean):

\[\large \int\limits \frac{ 1 }{ u^3 }du= \int\limits u^{-3}du\] you know how to integrate that right.

OpenStudy (christos):

to tell you the truth I already knew how to do it, I just think this exersise is wrong so I need to verify, look at the bottom http://screencast.com/t/XEJ9Dmj1S

OpenStudy (zzr0ck3r):

looks good

OpenStudy (christos):

my result: (1/2)*((-1/18) + (1/72))

OpenStudy (zzr0ck3r):

what is the problem?

OpenStudy (christos):

http://screencast.com/t/XEJ9Dmj1S this solution displays a different result

OpenStudy (christos):

Just making sure I didnt miss anything lol

OpenStudy (anonymous):

where did the \(\frac{1}{18}\) come from ?

OpenStudy (zzr0ck3r):

the solution looks fine

OpenStudy (christos):

3^2 = 9*2 = 18

OpenStudy (zzr0ck3r):

show us what you did, did u change your bounds when you substituted?

OpenStudy (zzr0ck3r):

3^2 = 3*3 = 9

OpenStudy (anonymous):

three squared is not equal to nine squared

OpenStudy (christos):

9*2(denominator) = 18

OpenStudy (zzr0ck3r):

18 = 2*3^2

OpenStudy (christos):

Right

OpenStudy (zzr0ck3r):

I m confused:)

OpenStudy (christos):

;p

OpenStudy (anonymous):

\[F(u)=-\frac{1}{4u^2}\] \[F(3)-F(6)=-\frac{1}{4}\left(\frac{1}{9}-\frac{1}{36}\right)\]

OpenStudy (christos):

where did the 1/4 come from

OpenStudy (christos):

there might be a probability that my solution and the other one are equal idk

OpenStudy (christos):

:D

OpenStudy (anonymous):

the u - sub gave a factor of \(\frac{1}{2}\) and the anti derivative gave another \(-\frac{1}{2}\)

OpenStudy (christos):

aah!

OpenStudy (christos):

hold on

OpenStudy (christos):

That was the trick, thank you

OpenStudy (anonymous):

yw